OpenStudy (anonymous):

Solve the equation. 2x − 3 + 4x = 21 x = −12 x = −6 x = 4 x = 6

OpenStudy (frozen15):

Simplifying 6x + 3 = 4x + 21 Reorder the terms: 3 + 6x = 4x + 21 Reorder the terms: 3 + 6x = 21 + 4x Solving 3 + 6x = 21 + 4x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-4x' to each side of the equation. 3 + 6x + -4x = 21 + 4x + -4x Combine like terms: 6x + -4x = 2x 3 + 2x = 21 + 4x + -4x Combine like terms: 4x + -4x = 0 3 + 2x = 21 + 0 3 + 2x = 21 Add '-3' to each side of the equation. 3 + -3 + 2x = 21 + -3 Combine like terms: 3 + -3 = 0 0 + 2x = 21 + -3 2x = 21 + -3 Combine like terms: 21 + -3 = 18 2x = 18

OpenStudy (anonymous):

1 Simplify 2x−3+4x to 6x−3 6x−3=21 2 Add 3 to both sides 6x=21+3 3 Simplify 21+3 to 24 6x=24 4 Divide both sides by 6 x=246 5 Simplify 246 to 4 x=4

OpenStudy (frozen15):

no fair , no one gets a medal , i don't like when someone (asker) doesn't give me a medal for helping .. even i do like helping